OpenStudy (anonymous):

Did I do this Algebra problem correctly? Simplify the expression. Write the answer using scientific notation 0.4(5.8x10^-3) I got, 2.32 x 10^-3 is this accurate? :D

OpenStudy (lgbasallote):

yes that is right

OpenStudy (anonymous):

Thank you! :D

OpenStudy (lgbasallote):

\[0.4 \times 5.8 \times 10^{-3} \implies (0.4 \times 5.8) \times 10^{-3} \implies 2.32 \times 10^{-3}\]